I tried multiple times and I can't get the right answers.

This is the exercise:

Nicole Limited is a company that produces machinery to customer orders, using a normal job-order cost system. It applies manufacturing overhead to production using a predetermined rate. This overhead rate is set at the beginning of each fiscal year by forecasting the year's overhead and relating it to direct labour costs. The budget for 2020 was as follows:

Direct labour

$1,803,000

Manufacturing overhead

901,500

As at the end of the year, two jobs were incomplete. These were 1768B, with total direct labour charges of $114,200, and 1819C, with total direct labour charges of $390,200. On these jobs, machine hours were 287 hours for 1768B and 647 hours for 1819C. Direct materials issued for 1768B amounted to $229,000, and for 1819C they amounted to $420,700.

Total charges to the Manufacturing Overhead Control account for the year were $898,500, and direct labour charges made to all jobs amounted to $1,577,800, representing 247,200 direct labour hours.

There were no beginning inventories. In addition to the ending work in process just described, the ending finished goods inventory account showed a balance of $562,520.

Sales for the year amounted to $6,205,300; cost of goods sold totalled $3,656,380; and sales, general, and administrative expenses were $1,856,800.

The above amounts for inventories and the cost of goods sold have not been adjusted for any over- or under-application of manufacturing overhead to production. It is the company's practice to allocate any over- or under-applied overhead to inventories and the cost of goods sold.

Q1 : Calculate the under- or over-applied manufacturing overhead for 2020.

Q2: Prorate the amount calculated in part (a) based on the ending balances (before prorating) of Work in Process, Finished Goods, and Cost of Goods Sold.

Allocation Cost of goods sold $1 Finished goods inventory WIP inventory Total %24 %24

Q3: Income statement for the company for the year. The income tax rate is 40%.

NICOLE LIMITED Income Statement 2$ %24
Answer #1

und UAAN 10 Oven applied dan ubacing uuuu 20 N Answer Page No 0 o undere come over applied manubacturing overhead for 20 20 BPage. No @ = (229000 + 420700) + (114200 + 390200) + |(114800 + 390200) X015] = 649700 + 504400 + 258200 = $ 1406300 ParticulPage No ③ Income statement Particular Sales. Adjusted cost of goods sold Gross porofit Operating expenses Met income before t